OpenStudy (anonymous):

For each pair of parametric equations below, eliminate the parameter and select the direct relationship between x and y. I almost have the correct answer, just not quite right!

OpenStudy (anonymous):

x=3-t, y=t^2-2t

OpenStudy (anonymous):

First you rearrange the x eqation so you get t by itself.